OpenStudy (anonymous):

solve for x in terms of y y=2^x-2^-x/2^x+2^-x

OpenStudy (anonymous):

help please

OpenStudy (anonymous):

\[ y=\frac{2^x-2^{-x}}{2^x+2^{-x}} \]

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

Do you know about \(\tanh(x)\)?

OpenStudy (anonymous):

\[ \tanh(x) = \frac{e^x-e^{-x}}{e^x+e^{-x}} \]

OpenStudy (anonymous):

no it is a part of logarithm

OpenStudy (anonymous):

\[ \tanh(x\ln2) = \frac{e^{x\ln 2}-e^{-x\ln 2}}{e^{x\ln 2}+e^{-x\ln2}} =\frac{2^x-2^{-x}}{2^x+2^{-x}} \]

OpenStudy (anonymous):

You want to do this with logarithm?

OpenStudy (anonymous):

yes not tangent

OpenStudy (anonymous):

Well if you have \[ \log_2(y)=\log_2(2^x-2^{-x})-\log_2(2^x+2^{-x}) \]How would you proceed?

OpenStudy (anonymous):

Hyperbolic tangent is the only way I know to do this.

OpenStudy (blockcolder):

Well, you can multiply numerator and denominator by 2^x: \[y=\frac{2^x(2^x-2^{-x})}{2^x(2^x+2^{-x})}=\frac{2^{2x}-1}{2^{2x}+1}\] then let u=2^(2x) so that \(y=\frac{u-1}{u+1}\). You then solve for u in terms of y, then put x back: \[y(u+1)=u-1\\ uy+y=u-1\\ uy-u=-1-y\\ u(y-1)=-(y+1)\\ u=-\frac{y+1}{y-1}\\ 2^{2x}=-\frac{y+1}{y-1}\] I hope you can take it from here. :)

OpenStudy (anonymous):

cross multiplication

OpenStudy (blockcolder):

No. Your goal was to write y in terms of x, right? And right now, x is already nearly alone. You just have to do something so that the x is not an exponent anymore. What is the inverse of exponentiation?

OpenStudy (anonymous):

i dont know exponentiation

OpenStudy (blockcolder):

Isn't your topic about logarithms?

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

the inverse of exponentiation is square root

OpenStudy (blockcolder):

No, I mean, say you have \(2^x=y\), it is true that \(x=\log_2(y)\), right?

OpenStudy (anonymous):

yeah right

OpenStudy (blockcolder):

So it's true that\[2x=\log_2{\left(-\frac{y+1}{y-1}\right)}\]right?

OpenStudy (anonymous):

it strue thats the final answer ?

OpenStudy (blockcolder):

Not yet. You have to divide both sides by 2 to finally get the final answer.

OpenStudy (anonymous):

how?@blockcolder

OpenStudy (anonymous):

@blockcolder

OpenStudy (blockcolder):

Just divide: \[\frac{2x}{2}=\frac{1}{2}\log_2{\left(-\frac{y+1}{y-1}\right)}\] So that you get \[x=\frac{1}{2}\log_2{\left(-\frac{y+1}{y-1}\right)}\]